OpenStudy (anonymous):

The following equation is already in factored form. Use the zero factor property to set the factors to 0 and solve. 7y(7y + 4)(8y + 7) = 0

OpenStudy (anonymous):

y=0 y=-4/7 y=-7/8

OpenStudy (saifoo.khan):

rite

OpenStudy (anonymous):

was i supposed to foil it>? thats what i did

OpenStudy (saifoo.khan):

Copy the answer, simply. lol

OpenStudy (anonymous):

so -4/7 is the smalest right?

OpenStudy (saifoo.khan):

no

OpenStudy (anonymous):

ohh i get what he did. just took 1 parenthesis at a time and set it to 0. so then -7/8 is the smallest?